Calculate the solubility in mol/dm of 40g of CuSO4 dissolved in 100g of water at 120 degrees Celsius

1
Expert's answer
2022-01-24T08:49:01-0500
  1. Solubility = number of moles of solution/vol of solution. 100g of water = 100 cm³ = 0.1 dm³.
  2. No of moles of CuSO4 in 40g = 40/160. = 0.25 mol.
  3. Solubility = 0.25/0.1. = 2.5 moldm−³.